# Boosting Adversarial Transferability for Hyperspectral Image Classification Using 3D Structure-invariant Transformation and Intermediate Feature Distance

### v1

@@ Abstract
Deep Neural Networks (DNNs) are vulnerable to adversarial attacks, which pose security challenges to hyperspectral image (HSI) classification technologies based on DNNs. In the domain of natural images, numerous transfer-based adversarial attack methods have been studied, among which input transformation is a common strategy to mitigate the overfitting of adversarial examples to surrogate models by applying diverse transformations to the images. Feature-space adversarial attacks, which modulate intermediate-layer representations of images, also demonstrate strong transferability. However, HSIs differ from natural images due to their high-dimensional and rich spectral information. Current research on HSI adversarial examples remains limited and faces challenges in fully utilizing the structural and feature information of images. To address these challenges, this paper proposes a novel method to enhance the transferability of the adversarial examples for HSI classification models. First, while keeping the image structure unchanged, the proposed method randomly divides the image into blocks in both spatial and spectral dimensions. Then, various transformations are applied on a block by block basis to increase input diversity and mitigate overfitting. Second, a feature distancing loss targeting intermediate layers is designed, which measures the distance between the amplified features of the original examples and the features of the adversarial examples as the primary loss, while the output layer prediction serves as the auxiliary loss. This guides the perturbation to disrupt the features of the true class in adversarial examples, effectively enhancing transferability. Extensive experiments demonstrate that the adversarial examples generated by the proposed method achieve effective transferability to black-box models on two public HSI datasets. Furthermore, the method maintains robust attack performance even under defense strategies.
@@
Hyperspectral image classification, Adversarial examples, Transferability, Feature distancing
@@ I Introduction
While obtaining geometric features of object surfaces through two-dimensional spatial imaging technology, hyperspectral images (HSIs) employ spectral demodulation techniques to interpret the total radiant energy of each pixel as a continuous spectrum across different bands [1]. Compared with natural images that have only three RGB channels, HSIs build a three-dimensional data cube that integrates the spatial-spectral information of ground targets, enabling the capture of more subtle differences between them. Due to this advantage, HSIs have important application value in meteorology [2], environmental protection [3], agriculture [4], and military fields [5].
@@
As a fundamental task in HSI applications, HSI classification aims to identify the substances represented by hyperspectral pixels. In recent years, DNNs have become the dominant approach for HSI classification due to their powerful feature extraction capabilities [6, 7, 8, 9]. Researchers have developed various methods based on convolutional neural networks (CNNs), recurrent neural networks (RNNs), and other DNN models to extract more discriminative spatial and spectral features from HSIs [10, 11, 12]. For example, Wang et al. [10] introduced a dynamic super-pixel normalization (DSN)-based DNN for HSI classification; Mahmoodi et al. [11] developed an effective method capable of focusing on both spectral and spatial features. When considering special circumstances such as insufficient training samples, more methods have been presented based on the techniques like few-shot learning for HSI classification [13, 14, 15].
Despite demonstrating significant advantages in HSI classification tasks, DNNs’ vulnerability to adversarial attacks and the reliability risks posed by adversarial examples still require focused attention. Many studies [16, 17] show that DNNs are highly vulnerable to adversarial attacks, which cause machine learning models to produce erroneous predictions via imperceptible subtle perturbations added to input data. These crafted inputs are referred to as adversarial examples. Current research on adversarial attacks mainly focuses on white-box settings where adversaries possess the target model’s internal information (including architecture, parameters, and training data) and leverage such knowledge to design targeted attacks. In contrast, black-box attacks do not require such access and are thus more realistic. Among the black-box attacks [18, 19], transfer-based attacks [20, 21] which utilize the transferability of adversarial examples to perform attacks are more flexible, where transferability refers to the property that adversarial examples trained on a known substitute model are also effective against structurally different target models.
@@
Researchers by now have paid much attention to adversarial examples in the natural image domain [22], with numerous methods proposed to enhance their transferability. These include classic approaches such as the FGSM method [23], and subsequent iterative methods [17, 24, 25]. However, adversarial examples generated using a substitute model often suffer from overfitting to the source model’s architecture and feature representations. To address this, various input transformation methods (e.g.,[26, 27, 28]) have been developed to improve transferability success rates. For example, the diverse input method (DI) [27] convert the input with a probability $p$, achieved by randomly resizing the image and padding zeros at its edges. Additionally, feature-space attack methods (e.g.,[21, 29, 30, 31]) aim to enhance transferability by modifying feature representations or altering the model’s attention, often guiding perturbation updates more purposefully through advanced loss functions. For example, Xu et al. [32] proposed a new black-box adversarial attack method called Mixup-Attack and its variant Mixcut-Attack, which finds common vulnerabilities between different networks by attacking the shallow features of a given substitute model.
However, due to the significant differences between HSIs and natural images, HSI classification needs to consider not only spatial but also spectral domain information, significantly increasing the complexity of generating perturbations. The data transformation-based attack methods proposed so far are generally more suitable for natural images. Performing transformations only in the spatial dimension fails to effectively utilize the redundant characteristics of HSI spectral data. Moreover, we argue that the structure of HSIs also contains a wealth of information, but previous methods (e.g., Shi et al. [33] approach utilizing superpixel segmentation and spectral clustering to enhance adversarial examples robustness and transferability) have not considered using this semantic information, leaving adversarial examples still suffer from the problem of overfitting to the substitute model. Furthermore, existing methods for generating adversarial examples usually rely on the classification layer information (e.g., using cross-entropy loss) to guide the generation of adversarial examples. However, the deep structure of neural networks leads to the problem of vanishing gradients during backpropagation [21], so relying solely on the distance between predicted labels as the optimization objective has significant limitations. Although methods like [32, 34] employ shallow-layer feature distances as losses, some models’ early layers may fail to learn truly salient features and semantic concepts [29], limiting the ability to guide perturbations in disrupting intra-image features and thus resulting in limited transferability.
Therefore, this study focuses on improving the transferability of adversarial examples in HSIs by enhancing input transformations and fully utilizing feature information. Wang et al. [35] observed in the natural image domain that more diverse image transformations can improve transferability while maintaining the structure unchanged. Given that HSIs typically contain abundant spectral bands (e.g.,the PaviaU dataset has 115 bands), we leverage this characteristic by dividing the input hyperspectral examples into blocks along both spatial and spectral dimensions and applying random transformations block by block to increase the diversity of the transformed images. Additionally, we perform attacks in the feature space to increase the distance between the original and adversarial examples. Inspired by the non-robust feature alignment method proposed by Zhu et al. [36] in the targeted attack approach, this paper designs a channel-weighted non-robust feature distancing loss (i.e., increasing inter-feature distances) from a non-targeted attack perspective. Initially, we calculate the weight of each channel to disrupt features that have a greater impact on model output. When implementing feature distance, we first enlarge the feature values of the original example, so that the feature values of the adversarial example become smaller than those of the original ones. This allows the perturbations to be updated in a direction that suppresses the original features within the adversarial examples. Finally, by integrating information from the model’s output layer, we construct a total loss function which forces the features of the adversarial example to deviate from the original distribution, thereby improving transferability.
@@
    We propose a novel attack framework to boost adversarial transferability for HSI classification by utilizing 3D structure-invariant transformations and intermediate feature distancing. While keeping the image structure unchanged, considering the numerous spectral bands of HSIs, we divided the image along both spatial and spectral dimensions. Subsequently, we applied random transformations to each block to enhance the diversity of the transformed images.
@@
    Previous methods mainly used classification layer information for loss calculation, resulting in insufficient utilization of image feature information. In this paper, we propose a new loss function that amplifies the features of the original examples and keeps them away from the features of adversarial examples. While calculating the information using channel weights, our method also integrates the classification information from the output layer to guide the exploration of more effective perturbations.
@@
    The transferability and defense experiments conducted on HSI datasets (PaviaU and HoustonU 2018) validate the effectiveness of the proposed method in generating adversarial examples with superior transferability and robustness against defenses.
The rest of this paper is organized as follows: Section 2 reviews related works relevant to this study. Section 3 provides a detailed explanation of the proposed adversarial attack methodology. Section 4 introduces the dataset information, experimental settings, and summarizes the experimental results. Section 5 concludes the study.
@@ II-A Traditional White-box Attack Methods
Gradient-based methods are a common and effective category of adversarial attack techniques. The core idea is to maximize the model’s loss using its gradient information to generate adversarial examples. Goodfellow et al. [23] proposed the Fast Gradient Sign Method (FGSM), which generates powerful adversarial examples based on the linear properties of CNNs. Wang et al. [17] and Madry et al. [24] further extended FGSM by decomposing the one-step perturbation generation into an iterative process, and proposed I-FGSM and PGD attacks. MI-FGSM [25] introduced a momentum term in iterative attacks, which accumulates previous gradient vectors to guide the calculation of the next gradient. Subsequent methods such as NI-FGSM [26] and VMI-FGSM [37] were also developed. Although these methods achieve high attack success rates on white-box models, their transferability to black-box models is often limited due to the significant differences in gradient information between models.
@@ II-B Input Transformation-based Attack Methods
Methods based on input transformations apply certain transformation enhancement operations to the input examples, increasing the diversity of examples fed into the model. This approach identifies model-discriminative regions common within the image, mitigating the overfitting of adversarial examples to the surrogate model. The DI [27] method is one representative approach. RDI [38] extends the DI approach by resizing the expanded image to its original size. TI [39], SIM [26], and Admix attack [28] are classical input transformation techniques that enhance diversity. Dong et al. [40]proposed a superpixel-guided attention attack method, where perturbations are no longer added at the pixel level, thereby improving the robustness of adversarial examples against defenses based on image processing. Long et al. [41] diversified inputs by performing transformations in the frequency domain. Wang et al. [42] addressed the limitations of existing input transformation methods ([26, 28]) in terms of adaptability and feature fusion effectiveness. Unlike previous methodologies, Guo et al. proposed MixCam [43], which enhances the transferability of adversarial examples from the perspective of shifting the regions that the model focuses on. Additionally, Wang et al. introduced BSR [44], improving the consistency of adversarial example attention heatmaps across different models. Liu et al. [45] designed a method with dynamically adjusted adaptive step sizes while enhancing input diversity, thereby further boosting the transferability of adversarial examples. These methods effectively improve attack transferability by increasing the diversity of inputs.
@@ II-C Advanced Loss Functions-based Attack Methods
Zhou et al. proposed Transferable Adversarial Perturbations (TAP) [21], which improve transferability by increasing the feature distance between the original and adversarial images in intermediate layers. Later, Ganeshan et al. introduced the Feature Disruptive Attack (FDA) [30], where the features of every layer in the model are disrupted. Inka et al. [31] modified the features of the original image to align its intermediate-layer representation becomes closer to that of another class. On the other hand, Wu et al. [46] approached the problem by extracting model attention, prioritizing the disruption of key features that may be adopted by multiple model architectures. Li et al. [47] introduced a method that mixes features of clean and adversarial examples, attenuating intermediate-level perturbations from clean features. Additionally, Wang et al. [48] proposed Diversifying the High-level Features to increases the diversity of features. Since the mapping distance of intermediate features correlates with the distance of the final output of neural networks [21], these methods increase the likelihood of erroneous predictions and enhance the success rate of black-box transfer attacks.
Figure 1: The overall workflow of the method. Before feeding the image into the substitute model, carefully crafted adversarial examples are partitioned in the spatial and spectral dimensions. Each block undergoes random transformations to enhance the diversity of the input. Subsequently, both adversarial and benign examples are input into the substitute model to extract feature maps from the intermediate layers. The weighted feature distancing loss and classification loss are computed to obtain the average gradient, which is iteratively updated to refine the perturbation. Here, $F_{x}^{l}$ and $F_{x_{adv}}^{l}$ denote the features of the benign and the adversarial examples at layer $l$, respectively.
II-D Adversarial Attacks and Defense in Remote Sensing
@@
Chen et al. [49] analyzed the vulnerability of CNN-based remote sensing image (RSI) recognition models to adversarial examples and found that models with different architectures exhibit varying degrees of vulnerability when trained on the same RSI dataset. Later, Xu et al. [50] and Chen et al. [51] investigated the performance of classic attack methods in remote sensing scene classification tasks. Additionally, Burnel et al. [52] used a Wasserstein Generative Adversarial Network-based approach to generate natural adversarial examples. Shi et al. [53] studied the vulnerabilities of multi-modal RSI classification networks and proposed a Multi-Feature Collaborative Adversarial Network. Further studies [33] explored the transferability of adversarial examples in HSIs by updating perturbations across spatial and spectral dimensions, these studies produced smoother adversarial examples while enhancing robustness against defense models. Previous studies that typically added perturbations globally, Liu et al. [54] achieved effective local pixel attacks, balancing attack effectiveness and cost. To mitigate the threat of adversarial attacks, various defenses have been proposed. Notably, Shi et al.[55] introduced an attention-based adversarial defense model that extracts and separates attack-invariant and attack-variant features from adversarial examples, using an attention-guided reconstruction loss to address the attention shift caused by perturbations. Xu et al. [56] leverage the global context information of HSIs to significantly improve model robustness against adversarial examples. They also proposed the Spatial-Spectral Self-Attention Learning Network [57], which integrates spatial and spectral information through a multi-scale spatial attention module and global spectral transformer. To the best of our knowledge, previous studies, with the except for [33], have not yet deeply explored how to improve the transferability of adversarial attacks in HSI classification, which we believe is also of significant research importance.

### v2

# Boosting Adversarial Transferability for Hyperspectral Image Classification Using 3D Structure-invariant Transformation and Weighted Intermediate Feature Divergence

@@ Abstract
Deep Neural Networks (DNNs) are vulnerable to adversarial attacks, which pose security challenges to hyperspectral image (HSI) classification based on DNNs. Studying adversarial attacks on DNN models for HSI classification helps reveal model vulnerabilities but also provides critical directions for improving their resilience. Numerous adversarial attack methods have been designed in the domain of natural images. However, different from natural images, HSIs contains high-dimensional rich spectral information, which presents new challenges for generating adversarial examples. Based on the specific characteristics of HSIs, this paper proposes a novel method to enhance the transferability of the adversarial examples for HSI classification using 3D structure-invariant transformation and weighted intermediate feature divergence. While keeping the HSIs structure invariant, the proposed method divides the image into blocks in both spatial and spectral dimensions. Then, various transformations are applied on each block to increase input diversity and mitigate the overfitting to substitute models. Moreover, a weighted intermediate feature divergence loss is also designed by leveraging the differences between the intermediate features of original and adversarial examples, which enhances the transferability of generated adversarial examples from the perspective of optimizing the perturbation generation process. It constrains the perturbation direction by enlarging the feature maps of the original examples, and assigns different weights to different feature channels to destroy the features that have a greater impact on HSI classification. Extensive experiments demonstrate that the adversarial examples generated by the proposed method achieve more effective adversarial transferability on three public HSI datasets. Furthermore, the method maintains robust attack performance even under defense strategies.
@@
Hyperspectral image classification, Adversarial examples, Adversarial transferability, Input transformation, Feature divergence
@@ I Introduction
While obtaining geometric features of object surfaces through two-dimensional spatial imaging technology, hyperspectral images (HSIs) employ spectral demodulation techniques to interpret the total radiant energy of each pixel as a continuous spectrum across different bands [1]. They build a three-dimensional data cube that integrates the spatial-spectral information of ground targets, enabling the capture of more subtle differences between them. Due to this advantage, HSIs have important application value in meteorology [2], environmental protection [3], agriculture [4], and military fields [5].
@@
In recent years, DNNs have become the dominant approach for HSI classification due to their powerful feature extraction capabilities [6, 7, 8, 9]. Researchers have developed various methods based on convolutional neural networks (CNNs), recurrent neural networks (RNNs), and other DNN models to extract more discriminative spatial and spectral features from HSIs [10, 11, 12]. For example, Wang et al. [10] introduced a dynamic super-pixel normalization (DSN)-based DNN for HSI classification; Mahmoodi et al. [11] developed an effective method capable of focusing on both spectral and spatial features. When considering special circumstances such as insufficient training samples, more methods have been presented based on the techniques like few-shot learning for HSI classification [13, 14, 15].
Despite demonstrating significant advantages in HSI classification tasks, DNNs are highly vulnerable to adversarial attacks [16, 17], which cause machine learning models to produce erroneous predictions via imperceptible subtle perturbations added to input example. These crafted inputs are referred to as adversarial examples. Therefore, it is of great significance to study adversarial attacks on the DNN models for HSI classification, which helps reveal the potential vulnerabilities of these models but also provides critical directions for improving their resilience.
@@
Previous research on adversarial attacks mainly focuses on white-box settings where adversaries possess the target model’s internal information such as architecture, parameters, and training data. Many classic methods have been proposed along this route, including FGSM [18] and subsequent iterative methods [17, 19, 20]. In contrast, black-box attacks do not require such access and generate adversarial examples using known substitute model, which are thus more realistic. Among the black-box attacks [21, 22], the attacks based on adversarial transferability [23, 24] are more flexible, where transferability refers to the fact that the adversarial examples obtained on the substitute model still have an attacking effect on the unknown target model. Nevertheless, adversarial examples generated using substitute model often overfit to the architecture and feature representations of that model. To address this, various input transformation methods (e.g.,[25, 26, 27]) have been developed to improve adversarial transferability. For example, the diverse input method (DI) [26] convert the input with a probability $p$, achieved by randomly resizing the image and padding zeros at its edges. At the same time, some methods enhance transferability by modifying feature representations or altering the model’s attention [24, 28, 29, 30]. For example, Xu et al. [31] proposed a new black-box adversarial attack method called Mixup-Attack, which finds common vulnerabilities between different networks by attacking the shallow-layer feature of a given substitute model.
However, since the HSIs contains not only a large amount of spatial information but also rich spectral information, this presents new challenges for generating adversarial examples. In previous methods, Shi et al. [32] focused on the robustness of generated adversarial examples against common image processing methods. They generated superpixel segmentation templates and spectral clustering templates in spatial and spectral dimensions, respectively, and constrained the perturbation values within each small block to be the same, improving the global smoothness of the generated perturbations. Although this approach takes into account the spectral information of HSIs, it has not explored the structure of HSIs which contains critical semantics information, and the issue of adversarial examples overfitting to the substitute model has not been alleviated. In addition, existing methods for generating adversarial examples for HSI classification models usually rely solely on classification layer information to guide perturbation optimization. Nevertheless, studies have found that adversarial perturbations generated from intermediate feature maps often have the highest transferability [33]. Although methods such as [31, 34] employ shallow-layer feature distance or feature approximation as loss functions, they failed to constrain the perturbation direction to weaken the features related to the true class. And in the feature space, they relied only on overall gradient information, without considering the different contributions of different channels in HSIs.
To address these issues, this study proposes a method to boost transferability of adversarial examples for HSI classification by using 3D structure-invariant transformation and weighted intermediate feature divergence. In the line of the transformation based black-box attacks, it exploits the large number of spectral bands of HSIs and partition HSIs into 3D blocks across both spatial and spectral dimensions. Random transformation strategies are applied to each block to obtain more diverse inputs, so as to alleviate the problem of overfitting to the substitute models. Meanwhile, inspired by Zhu et al. [33], a novel weighted intermediate feature divergence loss is also designed from the un-targeted attack perspective. To guide the perturbations to suppress features associated with the true class, it first enlarges the feature values of the original example by a fixed factor, ensuring that the adversarial example’s feature values are lower. Next, it computes the variance of each channel within the feature map to assign different weights to different channels, thereby destroying the features that have a greater impact on HSI classification.
@@
    By making full use of the large number of spectral bands in HSIs, we propose a 3D structure-invariant transformation method to divide HSIs into blocks in both spatial and spectral directions and apply random transformation strategies to each block. While keeping the global structure unchanged and maintain global semantic information, it enhances the diversity of HSIs and alleviates the problem of generated adversarial examples overfitting to the substitute model.
@@
    We design a weighted intermediate feature divergence loss that enhances the transferability of generated adversarial examples by increasing the differences between the intermediate features of the original and adversarial examples. It constrains the perturbation direction by enlarging the feature maps of the original example, and assigns different weights to different feature channels to destroy the features that have a greater impact on HSI classification.
@@
    We propose a novel attack framework to boost the transferability of adversarial examples for HSI classification by utilizing 3D structure-invariant transformations together with the weighted feature divergence loss. Our framework is combined with two gradient-based attack methods. Extensive experiments on three real HSIs datasets show that the proposed method outperforms the baseline approaches with significant advantages, demonstrating its effectiveness.
The rest of this paper is organized as follows: Section 2 reviews related works relevant to this study. Section 3 provides a detailed explanation of the proposed adversarial attack methodology. Section 4 introduces the datasets information, experimental settings, and summarizes the experimental results. Section 5 concludes the study.
@@ II-A Traditional White-box Attack Methods
Gradient-based methods are a common and effective category of adversarial attack techniques. The core idea is to maximize the model’s loss using its gradient information to generate adversarial examples. Goodfellow et al. [18] proposed the Fast Gradient Sign Method (FGSM), which generates powerful adversarial examples based on the linear properties of CNNs. Wang et al. [17] and Madry et al. [19] further extended FGSM by decomposing the one-step perturbation generation into an iterative process, and proposed I-FGSM and PGD attacks. MI-FGSM [20] introduced a momentum term in iterative attacks, which accumulated previous gradient vectors to guide the calculation of the next gradient. Subsequent methods such as NI-FGSM [25] and VMI-FGSM [35] were also developed. Although these methods achieve high attack success rates on white-box models, their transferability to black-box models is often limited due to the significant differences in gradient information between models.
@@ II-B Input Transformation-based Attack Methods
The approach based on input transformation enhances the diversity of input data by applying specific data transformation operations to the inputs. The DI [26] method is one representative approach. TI [36], SIM [25], and Admix attack [27] are classical input transformation techniques that enhance diversity. Dong et al. [37] proposed a superpixel-guided attention attack method, which constrains the perturbation values within each superpixel to be the same, thereby improving the robustness of adversarial examples against defenses based on image processing. Long et al. [38] diversified inputs by performing transformations in the frequency domain. Unlike previous methodologies, Guo et al. proposed MixCam [39], which enhances adversarial transferability by shifting the regions of model attention. Additionally, Wang et al. introduced BSR [40], improving the consistency of adversarial example attention heatmaps across different models. Liu et al. [41] proposed a local mixup method to improve the existing mixup method. The transformed examples obtained by these methods can be extended the model’s attention areas, mitigate the overfitting to the substitute model during the generation of adversarial examples, and thereby improve the success rate of the attack.
@@ II-C Advanced Loss Functions-based Attack Methods
Yosinski et al. [28] were the first to propose that the features extracted by the network have transferability across models with different architectures. Later, Zhou et al. proposed Transferable Adversarial Perturbations (TAP) [24], which improve transferability by increasing the feature distance between the original and adversarial images in intermediate layers. Inka et al. [30] modified the features of the original image to align its intermediate-layer representation becomes closer to that of another class. On the other hand, Wu et al. [42] approached the problem by extracting model attention, prioritizing the disruption of key features that may be adopted by multiple model architectures. Li et al. [43] introduced a method that mixes features of clean and adversarial examples, attenuating intermediate-level perturbations from clean features. Additionally, Wang et al. [44] also proposed to increases feature diversity by diversifying the high-level features. Since the mapping distance of intermediate feature maps correlates with the distance of the final output of neural networks [24], these methods increase the likelihood of erroneous predictions and enhance the success rate of transferability attacks.
II-D Adversarial Attack methods in Remote Sensing
In recent years, research in the area of adversarial examples within the remote sensing field has been steadily increasing. Chen et al. [45] analyzed the vulnerability of CNN-based remote sensing image (RSI) recognition models to adversarial examples and found that models with different architectures exhibit varying degrees of vulnerability when trained on the same RSI dataset. Later, Xu et al. [46] and Chen et al. [47] investigated the performance of classic attack methods in remote sensing scene classification tasks. Burnel et al. [48] used a Wasserstein Generative Adversarial Network-based approach to generate more natural adversarial examples for remote sensing. Methods based on implicit neural representations and feature approximation (i.e., approximating the features of the original image to a virtual image not belonging to any class) provided new insights for this research [49, 34]. The latest DMFAA method [50] is based on training a substitute model using a distillation model, integrating features from other models. Shi et al. [32] obtained smoother examples in the spatial and spectral dimensions through constraints based on superpixels and clustering, exploring their adversarial transferability and robustness. Previous studies that typically added perturbations globally, Peng et al. [51] proposed a semi-black-box adversarial attack method based on one-pixel attacks.
@@ II-D Adversarial Attack methods in Remote Sensing
To mitigate the threat of adversarial attacks, various defenses have been proposed. Shi et al.[52] introduced an attention-based adversarial defense model that extracts and separates attack-invariant and attack-variant features from adversarial examples, using an attention-guided reconstruction loss to address the attention shift caused by perturbations. Xu et al. [53] leverage the global context information of HSIs to significantly improve model robustness against adversarial examples. They also proposed the Spatial-Spectral Self-Attention Learning Network [54], which integrates spatial and spectral information through a multi-scale spatial attention module and global spectral transformer. Similarly, [55, 56] integrated spatial and spectral information for defense, indicating that holistic consideration of global information is an effective defense strategy. Furthermore, Zhao et al. [57] proposed a novel anti-perturbation network that enhances the robustness of HSI classification under adversarial attacks by integrating pixel-level features and superpixel priors. To the best of our knowledge, previous studies have paid little attention to improving the transferability of adversarial attacks in HSI classification, which we believe is also of significant research importance.